# ConfuGuard: A Low False-Positive System for Detecting Malicious and Stealthy Typosquatting Threats in Package Registries

### v2

@@ 5.5.2 Approach
Our analysis of false-positive data (section 4.2) showed many reasons for which legitimate packages may have names resembling other trusted resources. We iteratively developed 10 rules that utilize metadata features (e.g., version history, maintainers) to distinguish package confusion attacks from benign packages. We apply these rules to assess the benignity of fetched neighbors from section 5.4 and reduce the false positive rate. We added several additional rules during deployment. During implementation, we identify rules that require description analysis (e.g., R1, R2, R3) or additional knowledge (e.g., R8). To handle these cases, we use an LLM to analyze each package and generate the corresponding rules. We provide the package description and relevant metadata as input, refining the prompt iteratively based on feedback from security analysts. The final prompts are available in our artifact (Data Availability and Research Ethics).
Table 3 summarizes the goals and implementation details for each rule. Once our name-based detector flags a package as suspicious, we retrieve its metadata and apply benignity check rules, using a simple weighted sum to calculate a risk score. We have explored both manually-assigned weights and learning them from regression. The simple scoring function facilitates feedback from analysts.
@@
Use in Production
ConfuGuard is integrated into our industry partner’s security scanning system for software packages. They had previously relied primarily on content-based malware scanners, but these are so computationally costly that they cannot be applied at scale. Their former approach for scalable detection of confusion attacks was a lexical check (Levenshtein distance) which had an unacceptably high false positive rate. ConfuGuard replaced that approach and complements the malware scanners. A team of analysts triages alerts from ConfuGuard and the other scanners to make a threat assessment. If they deem a package to be malicious, that information is propagated to our partner’s customers through their security feeds.
Experience-Driven Optimizations
We trialed several versions and parameterizations of ConfuGuard over several months to enhance its performance. We share four examples. First, to improve accuracy across ecosystems, we tailored the neighbor search query (Part 4). For instance, lengthy identifiers, e.g., Maven’s artifact_id and Golang’s domain name, diminish the effectiveness of embedding similarity. To address this, we opted to compute separate similarity measurements for each component of the package names and apply an edit distance threshold. If the difference exceeds two-thirds of the original name’s length, it is unlikely to confuse users. Second, within our threat model, we determined that changes to both identifiers in Maven, Golang, and Hugging Face were unlikely to confuse users and no such attacks were observed. Consequently, we excluded these from our algorithm. Third, we refined our metadata checks for benignity. We added rules R11, R12, and R13 based on observed patterns. Fourth, for performance, we experimented with quantized versions of the embedding model. Ultimately we found the latency gain a poor bargain for the resulting accuracy degradation.
@@ 6 Evaluation
To evaluate ConfuGuard, we pose five Evaluation Questions (EQs) to assess its performance at the component level (EQ1EQ3) and the system level (EQ4EQ5). At the component level, we measure the effectiveness of novel mechanisms introduced in ConfuGuard. At the system level, we evaluate its integrated functionality and scalability compared to baseline tools (section 6.1), and ability to detect real-world package confusion threats. Additionally, we compare our approach to SOTA methods to benchmark its effectiveness. An overview of the evaluation process is illustrated in Figure 3.
Component-level
We assess how individual components contribute effectively to the overall system.
  • EQ1: Performance of Embedding Model. What is the accuracy and efficiency of our embedding model? (section 5.2)
  • EQ2: Neighbor Search Accuracy. How effective is the neighbor search approach? (section 5.4)
  • EQ3: Metadata Verification Accuracy. How much does the benignity filter reduce false positive rates? (section 5.5)
System-level
We examine the performance of the full ConfuGuard system and compare to other approaches.
@@ 6.1 Baseline and Evaluation Datasets
State-of-the-Art Baselines
We compare our system to the Levenshtein distance approach [63], and Typomind [39]. We consider OSSGadget [36] out of scope because it only handles lexical confusions and its latency for long package names makes it unsuitable for production. These are all state-of-the-art open-source tools.
Evaluation Datasets
We evaluate using two datasets:
  • NeupaneDB: 1,840 packages from [39], including 1,239 confirmed real-world package confusion attacks, and 601 manually labeled data we analyzed in section 4.2.
  • ConfuDB: 1,561 packages triaged by security analysts, collected during the development and refinement of ConfuGuard (section 5.6).
6.2 EQ1: Perf. of Embedding Model
@@
6.2.1 Effectiveness
Method
We evaluate the effectiveness of embedding-based similarity detection by comparing three approaches:
@@ 6.2.1 Effectiveness
    Levenshtein-Distance, calculates the number of single-character edits required to change one package name to another.
@@
    Pre-trained FastText [14] (cc.en.300.bin), used in the SOTA work Typomind [39], employs the general-purpose embedding model cc.en.300.bin [14] to capture semantic relationships.
@@
    Fine-tuned FastText (Ours), which we have adapted using a corpus specifically composed of package names to better capture domain-specific similarities.
To systematically compare these methods, we construct a balanced test set consisting of both positive and negative pairs, with each category containing 1,239 data points.
  • Positive Pairs: 1,239 real package confusions from NeupaneDB [39].
  • Negative Pairs: Created by randomly pairing unrelated package names from registries, ensuring low similarity scores across all tested methods. These pairs are guaranteed not to represent package confusion relationships.
[…]
@@
Figure 4: Performance Metrics of the Fine-Tuned Model: (a) Threshold Accuracy and (b) ROC Curve.
Result
Table 4 presents the comparison between our fine-tuned model and baseline approaches. The results show that our approach outperforms both the Levenshtein Distance and Pretrained models, achieving the highest F1 score for positive pairs (0.95) while maintaining strong performance on negative pairs (0.98).
Figure 4 illustrates the performance metrics of our approach through a grid search, highlighting the relationship between similarity threshold and accuracy, as well as the ROC curve for the fine-tuned model. The optimal similarity threshold of 0.93, identified through this search, is used in the production system. With an AUC of 0.94, our model demonstrates a balanced trade-off between precision and recall for both positive and negative pairs.
6.2.2 Efficiency
Method
To assess the efficiency of embedding database creation, we evaluated the overhead of embedding database creation step (section 5.3). We evaluated the embedding database creation efficiency by measuring the throughput, latency, and overall overhead associated.
Table 5: Evaluation of embedding model efficiency, HNSW indexing overhead, and total embedding database creation overhead.
Ecosystem Throughput Latency Indexing Time Total Overhead
(pkgs/s) (ms) (s) (s)
NPM 7705.99 0.13 7.42 650.19
PyPI 13581.15 0.07 4.39 46.85
RubyGems 8694.68 0.12 4.07 28.18
Maven 4887.19 0.20 4.91 134.47
Golang 8490.36 0.12 7.80 232.93
Hugging Face 5027.95 0.20 4.46 166.67
Result
Table 5 presents the efficiency measurement. Overall, the quantized models demonstrate strong performance with minimal overhead. The HNSW indexing adds negligible processing time, requiring less than 10 seconds additional overhead per table.
@@ 6.3 EQ2: Neighbor Search Accuracy
Methods
We use a threshold of 0.93 obtained from EQ1. To evaluate neighbor search performance, we analyzed suspicious packages identified by Typomind.
Results
Our neighbor search algorithm accurately detected 99% of the 1,239 real-world package confusion attacks from NeupaneDB [39]. Moreover, our method effectively flagged impersonation squatting attacks targeting hierarchical package names, notably identifying reported cases on Hugging Face [43] and Golang [16] that earlier methods overlooked. These findings confirm that our neighbor search algorithm achieves state-of-the-art performance.
@@ 6.4 EQ3: Benignity Check Accuracy
Method
Starting with the raw embedding output (i.e., packages flagged for name-based suspicion), we apply the benignity check described in section 5.5.
We first use cross-validation to learn the weights and measure false positive rates (FPR) using NeupaneDB. We then computed SHAP values by learning the weighed sum through regression to quantify the importance of each metric defined in Table 2 [33]. The resulting SHAP plot provides clear insights into how these metrics contribute to our system’s decision-making process. Additionally, we use regression to learn the weights
Table 6: Accuracy metrics for ConfuGuard and Typomind, detailing true/false positives and negatives, as well as recall, precision, F1 score, and accuracy for active, stealthy, and benign threats. Although Typomind flags all packages as suspicious, our benignity filter substantially reduces false positives in production. Due to availability of metadata, we evaluated ConfuGuard on 1,043 packages from NeupaneDB and 1,202 ConfuDB. We note that ConfuGuard exhibits lower benign classification accuracy in ConfuDB because it flags more complex cases (e.g., compound squatting) as potential threats, a nuance that Typomind does not capture.
Dataset Total Threat Type Sub-total ConfuGuard Typomind
@@
Active 1,239 386 0 0 57 0.87 1.00 0.93 0.87 1000 0 0 239 0.80 1.00 0.89 0.81
NeupaneDB 1,840 Stealthy 137 80 0 0 56 0.59 1.00 0.74 0.59 137 0 0 0 1.00 1.00 1.00 1.00
Benign 464 0 72 391 0 0.00 0.00 0.00 0.84 0 464 0 0 0.00 0.00 0.00 0.00
Active 71 41 0 0 4 0.91 1.00 0.95 0.91 40 0 0 31 0.56 1.00 0.72 0.56
ConfuDB 1,561 Stealthy 230 110 0 0 24 0.82 1.00 0.90 0.82 111 0 0 113 0.50 1.00 0.66 0.50
Benign 1,260 0 532 491 0 0.00 0.00 0.00 0.48 0 322 612 0 0.00 0.00 0.00 0.65
Result
Figure 5 shows the SHAP value plot which indicates that our rules used in the benignity filter are useful. The plot highlights that distinct purpose (R2) and suspicious intent (R10) are the most influential features driving the model’s predictions, suggesting that packages showing suspicious or unusual behavior strongly push the model toward a malicious classification. Conversely, features such as known maintainers (R8) and fork packages (R3) generally pull the model’s output toward benign, indicating they lower the likelihood of a threat. Meanwhile, no clear description also shows a notable positive effect, aligning with the idea that missing documentation can be an indicator of malicious intent (R9).
We use cross-validation to learn parameters and estimate the trade-off between reducing false positives and increasing false negatives. Our benignity filter correctly flagged 399 out of 464 (86%) false positives in the human labeled set. In the real-world attack set, it failed to classify 53 true positives out of 443.333Only 443 out of 1,239 real attacks still have available metadata for our analysis. A detailed investigation revealed two main causes: (1) missing metadata and (2) attacks that had been removed, with the package now maintained by npm, which the filter considers a valid maintainer.
These results confirm that supplementary heuristics beyond raw name similarity reduce false positives while retaining high recall for genuinely malicious confusions.
Figure 5: SHAP value plot for metadata verification using learnable features (i.e., excluding R13, R14) from Table 3. The x-axis shows each feature’s impact on model output (positive values are towards benginity). Feature values are color-coded: red indicates high values with a strong, direct impact on the model output.
@@ 6.5 EQ4: Discovery of New Package Confusions
Method
To evaluate the effectiveness of our package confusion detection system, we deployed ConfuGuard in a production environment for three month. During this period, flagged packages were analyzed using a commercial malware scanner and reviewed by threat analysts for detailed insights.
Result
Table 6 presents the package confusion attack that ConfuGuard identified during three months deployment period (i.e., ConfuDB). Within this database, 1,260 packages were labeled as false positives. [^4] The high false positive rate in ConfuDB is primarily due to the difficulty of confirming stealthy confusion attacks, which were left unreviewed by analysts. Among the remaining 301 suspicious packages, 230 (76%) exhibited stealthy confusion attack behavior, 63 (21%) were identified as malware, 6 (2%) were categorized as vulnerabilities, and 2 (0.7%) fell into other anomaly categories.

### v3

@@ 5.5.2 Approach
Our analysis of false-positive data (section˜4.2) showed many reasons why legitimate packages may have names resembling other trusted resources. Initially, we iteratively developed 11 rules that utilize metadata features (e.g., version history, maintainers) to distinguish package confusion attacks from benign packages. We apply these rules to assess the benignity of package neighbors (section˜5.4) in order to reduce the false positive rate. We added several additional rules during deployment. Table˜3 summarizes the goals and implementation for each rule. Once our name-based detector flags a package we retrieve its metadata and apply benignity check rules. A risk score is calculated using a simple weighted sum, which facilitated analyst feedback compared to a learned approach.
{listing}
[ht]
# ------------ system_prompt (excerpt) ------------
You are an AI assistant helping to identify false positives in potential typosquat packages.
Analyze the given package names and descriptions to determine if a package is a legitimate alternative or a malicious typosquat.
@@
[…]
- TRUE if: Names are confusingly similar.
- FALSE if: Names are clearly distinct or branded
[…]
- TRUE if: High edit distance AND distinct purpose AND good docs
- FALSE if: Low edit distance OR similar purpose OR poor docs
[…]
# ------------ user_prompt (excerpt) --------------
[…]
Analyze the information above and answer the questions below using TRUE or FALSE.
[…]
Excerpts of the system and user prompts used in the benignity filter. Full prompts are in our artifact (Data Availability). We refined both prompts iteratively with feedback from security analysts, ensuring that the LLM’s analysis aligns with the rule set in Table˜3.
To realize this filter, we first distinguished the rules that can be assessed using Boolean algebra over package metadata (e.g., R4), from those that require extensive analysis (e.g., R2) or subjective reasoning (e.g., R9). The former we implemented with conventional software, and for the latter we used an LLM. We tried two prompting approaches: rule-based and “LLM-as-judge” [27]. For rule-based, we provided detailed rules imitating security analysts’ reasoning (e.g., specific metadata they referenced). This detected known attacks (accuracy of 0.82 on NeupaneDB), but overfit, raising false positives on new data (accuracy of 0.53 on ConfuDB as of March 2025). We also tried the “LLM-as-a-Judge” pattern, casting the model as an evaluator guided by a structured decision framework but without excessive detail [27]. This pattern performed better, with accuracy of 0.83 on NeupaneDB and 0.68 on the latest ConfuDB (Table˜5).
Section˜5.5.2 shows the final system and user prompts. We used OpenAI’s o4-mini in our implementation.
[…]
Use in Production ConfuGuard is integrated into our industry partner’s security scanning system for software packages. They had previously relied primarily on content-based malware scanners, but these are too computationally costly to scale well. The partner’s former approach for first-pass detection of confusion attacks was a lexical check (Levenshtein distance) which had a high false positive rate. ConfuGuard replaced that approach. Analysts triage alerts from ConfuGuard and the other scanners to make a threat assessment. If they deem a package to be malicious, that information is propagated to our partner’s customers via security feeds.
Experience-Driven Optimizations During deployment we made several optimizations. We share three examples. First, we enhanced our metadata checks by analyzing false positives and adding rules R11R15 to better distinguish benign packages (Table˜3). Second, we improved accuracy by splitting long identifiers, such as Golang’s domain-based names, into their constituent parts and computing similarity on each part. Third, we found the LLM-as-Judge pattern was more effective than detailed prompts (section˜5.5).
@@ 6 Evaluation
To evaluate ConfuGuard, we pose five Evaluation Questions (EQs) to assess its performance at the component level (EQ1EQ3) and the system level (EQ4EQ5). At the component level, we measure the effectiveness of novel mechanisms introduced in ConfuGuard. At the system level, we evaluate its integrated functionality and scalability compared to baseline tools (section˜6.1), and ability to detect real-world package confusion threats. Additionally, we compare our approach to SOTA methods to benchmark its effectiveness. An overview of the evaluation process is illustrated in Figure˜3.
Component-level We assess how individual components contribute effectively to the overall system.
  • EQ1: Performance of Embedding Model. What is the accuracy and efficiency of our embedding model? (section˜5.3)
  • EQ2: Neighbor Search Accuracy. How effective is the neighbor search approach? (section˜5.4)
  • EQ3: Benignity Check Accuracy. How much does the benignity filter reduce false positive rates? (section˜5.5)
System-level We examine the performance of the full ConfuGuard system and compare to other approaches.
@@ 6.1 Baseline and Evaluation Datasets
State-of-the-Art Baselines We compare our system to the Levenshtein distance approach [75], and Typomind [49]. We consider OSSGadget [46] out of scope because it only handles lexical confusions and its latency for long package names makes it unsuitable for production. These are all state-of-the-art open-source tools.
Evaluation Datasets We evaluate using two datasets:
  • NeupaneDB: 1,840 packages from [49], including 1,239 confirmed real-world package confusion attacks, and 601 manually labeled data we analyzed in section˜4.2.
  • ConfuDB: 2,361 packages triaged by security analysts, collected during the development and refinement of ConfuGuard (section˜5.6).
6.2 EQ1: Performance of Embedding Model
@@
6.2.1 Effectiveness of Embedding Model
Method We evaluate the effectiveness of embedding-based similarity detection by comparing three approaches:
@@ 6.2.1 Effectiveness of Embedding Model
    Levenshtein Distance, calculating the number of single-character edits required to change one package name to another.
@@
    Pre-trained FastText [9] (cc.en.300.bin), as used in the SOTA work Typomind to capture semantic relationships [49].
@@
    Fine-tuned FastText (Ours), trained with a corpus of package names to better capture domain-specific similarities.
To systematically compare these methods, we construct a balanced test set consisting of both positive and negative pairs, with each category containing 1,239 data points [8, 19].
  • Positive Pairs: All 1,239 of the real package confusions from NeupaneDB, which were confirmed and removed from SPRs [49].
  • Negative Pairs: Created synthetically by randomly pairing unrelated package names from registries, ensuring low similarity scores across all tested methods. These pairs are guaranteed not to represent package confusion relationships.
For each approach, we applied a predefined similarity threshold to classify package pairs as potential package confusion attacks. The threshold was selected via a grid search over values from 0 to 1 in 0.01 increments to optimize Precision and Recall, ensuring effective identification of true confusions. Pairs with similarity scores above the threshold were classified as positive, while those below were classified as negative. False positives were subsequently filtered using our false-positive verification process (section˜5.5). We compare the Precision, Recall, and F1 scores for each approach.
Result Table˜4 presents the comparison between our fine-tuned model and baseline approaches. The results show that our approach outperforms both the Levenshtein Distance and Pretrained models, achieving the highest F1 score for positive pairs (0.95) while maintaining strong performance on negative pairs (0.98).
@@
We performed a grid search and analyzed the ROC curve to assess how the similarity threshold affects accuracy and to plot the ROC curve of our fine-tuned model. This analysis identified 0.93 as the optimal threshold, which we adopted in production. With an AUC of 0.94, the model demonstrates a balanced trade-off between precision and recall for both positive and negative pairs.
6.2.2 Efficiency of Embedding Model
Method To assess the efficiency of embedding database creation, we evaluated the cost of creating the embedding database (section˜5.3). We measured the throughput, latency, and overall overhead associated.
Result Creating the embedding database can be done in a reasonable amount of time, even for large SPRs. Total database preparation cost varied from 28–650 seconds (RubyGems, NPM), largely driven by the size of the dataset in each ecosystem. Throughputs ranged from $\sim$4.9–13.6 K pkgs/sec (Maven, PyPI), with per-package latencies consistently below 0.2 ms. HNSW indexing adds $<$ 8 seconds of upfront processing for each ecosystem.
@@ 6.3 EQ2: Neighbor Search Accuracy
Method We use a threshold of 0.93 obtained from EQ1. To evaluate neighbor search performance, we analyzed suspicious packages identified by Typomind.
Result Our neighbor search algorithm accurately detected 99% of the 1,239 real-world package confusion attacks from NeupaneDB [49]. Moreover, our method effectively flagged impersonation squatting attacks targeting hierarchical package names, notably identifying reported cases on Hugging Face [54] and Golang [11] that earlier methods overlooked. These findings confirm that our neighbor search algorithm achieves state-of-the-art performance.
@@ 6.4 EQ3: Benignity Check Accuracy
Method Using the raw embedding output (i.e., packages flagged for name-based suspicion), we apply the benignity check from section˜5.5.
We first use cross-validation to learn the weights and measure false positive rates (FPR) using NeupaneDB. We then computed SHAP values [43] by learning the weighed sum through regression to quantify the importance of each metric defined in Table˜2. The resulting SHAP plot provides clear insights into how these metrics contribute to our system’s decision-making process. Additionally, we use regression to learn the weights.
Table 5: Accuracy metrics for ConfuGuard and Typomind, detailing true/false positives and negatives, as well as recall, precision, F1 score, and accuracy for active, stealthy, and benign threats. Although Typomind flags all packages as suspicious, our benignity filter substantially reduces false positives in production. Due to availability of metadata, we evaluated ConfuGuard on 1,819 packages from NeupaneDB and 2,034 from ConfuDB. We note that ConfuGuard exhibits lower benign classification accuracy in ConfuDB because it flags more complex cases (e.g., compound squatting) as potential threats, a nuance that Typomind does not capture. ConfuDB data was collected between January and June 2025.
Dataset Total Threat Type Sub-total ConfuGuard (Ours) Typomind
@@
Active 1,239 1,103 0 0 131 0.89 1.00 0.94 0.89 1,000 0 0 239 0.80 1.00 0.89 0.81
NeupaneDB 1,840 Stealthy 121 68 0 0 53 0.56 1.00 0.72 0.56 121 0 0 0 1.00 1.00 1.00 1.00
Benign 480 0 117 347 0 0.00 0.00 0.00 0.72 0 480 0 0 0.00 0.00 0.00 0.00
Overall 1,840 1,171 117 347 184 0.86 0.90 0.88 0.83 1,121 480 0 239 0.82 0.70 0.76 0.61
Active 133 61 0 0 72 0.53 1.00 0.70 0.53 71 0 0 62 0.53 1.00 0.72 0.56
ConfuDB 2,034 Stealthy 409 238 0 0 171 0.74 1.00 0.85 0.73 219 0 0 190 0.64 1.00 0.78 0.64
Benign 1,492 0 414 1078 0 0.00 0.00 0.00 0.72 0 560 932 0 0.00 0.00 0.00 0.61
Overall 2,034 299 414 1078 243 0.55 0.42 0.48 0.68 290 560 932 252 0.62 0.28 0.39 0.62
Result Figure˜4 shows the SHAP value plot which indicates that our rules used in the benignity filter are useful. The plot highlights that distinct purpose (R2) and suspicious intent (R10) are the most influential features driving the model’s predictions, suggesting that packages showing suspicious or unusual behavior strongly push the model toward a malicious classification. Conversely, features such as known maintainers (R8) and fork packages (R3) generally pull the model’s output toward benign, indicating they lower the likelihood of a threat. Meanwhile, no clear description also shows a notable positive effect, aligning with the idea that missing documentation can be an indicator of malicious intent (R9).
Figure 4: SHAP value plot for metadata verification over the learnable features from Table˜3 (i.e., excluding R13, R14, R15). The x-axis shows each feature’s impact on model output, where positive values are towards benignity. Feature values are color-coded: red indicates high values with a strong, direct impact on the model output.
We use cross-validation to learn parameters and estimate the trade-off between reducing false positives and increasing false negatives. Our benignity filter correctly flagged 347 out of 480 (72%) false positives in the human labeled set. In the real-world attack set, it failed to classify 131 true positives (10%) out of the 1,234 that had available metadata. A detailed investigation revealed two main causes: (1) missing metadata, and (2) attacks that had been removed, with the package now maintained by npm, which the filter considers a valid maintainer.
These results confirm that by including supplementary heuristics beyond raw name similarity, false positive rates can be reduced while retaining high recall for genuinely malicious confusions.
@@ 6.5 EQ4: Discovery of New Package Confusions
Method To evaluate the effectiveness of our package confusion detection system, we deployed ConfuGuard in a production environment for three month. During this period, flagged packages were analyzed using a commercial malware scanner and reviewed by threat analysts for detailed insights.
Result Table˜5 summarizes the package confusion attacks identified by ConfuGuard during the January–June 2025 deployment period (i.e., ConfuDB). Within this database, 1,686 packages were labeled as false positives. [^3] The high false positive rate in ConfuDB is primarily due to the difficulty of confirming stealthy confusion attacks, which were left unreviewed by analysts. Among the 630 remaining suspicious packages, 119 (19%) were identified as executable malware, 7 (1%) launched HTTP-flood or DDoS attacks, 6 (1%) loaded external scripts from unknown URLs or CDNs, and 1 (<1%) harvested analytics data and sent it to the attackers, and 497 (79%) carried out stealthy confusion attacks without obvious malicious content.
Table˜6 shows the different kinds of package confusion threats identified by ConfuGuard during its 6 months of production use. ConfuGuard effectively detects advanced threats (section˜4.1) — including compound and impersonation squatting — across platforms such as Maven, Golang, and Hugging Face.
@@
Table 6: ConfuGuard’s detection counts by confusion-attack category in production. This total exceeds the numbers in Table˜5, as it reflects the cumulative counts collected over the six-month production deployment. Traditional lexical confusion types (targeted by Typomind and SpellBound) are listed first, then additional semantic confusions escaping prior work, and finally new patterns (Table˜2).
Attack Category Counts (Pct.) Typomind
Traditional Lexical Confusion 431 (68.4%) 431
Additional Semantic Confusion 69 (11.0%) 0
Impersonation Squatting 78 (12.4%) 0
Compound Squatting 50 ( 7.9%) 0
Domain Confusion 2 ( 0.3%) 0
Total 630 (100%) 431 (68%)